OpenStudy (anonymous):

I have this problem...help. 30m/= ft/s?

OpenStudy (anonymous):

30 m/h = ft/s?

OpenStudy (anonymous):

jliberary is closing....got to go.....my emial is etina02@yahoo.com.....

OpenStudy (anonymous):

How do you convert 30m/h = ft/s?

OpenStudy (btaylor):

30 mph x (1 hour / 3600 sec) --> .00833333 mi/sec .00833333 mi/sec x (5280 ft / 1 mi) --> 44 ft/sec

OpenStudy (anonymous):

2.7340E-2ft/s

OpenStudy (anonymous):

It seems unlikely that speed is measured in meters per hour, so I will convert from miles per hour.\[30mph=\frac{30 miles}{1hour}\frac{5280ft}{1mile}\frac{1hr}{3600\sec}=44fps\]

OpenStudy (anonymous):

m isn't meters?

OpenStudy (anonymous):

Usually is, but it doesn't seem likely that if a speed was originally in meters per hour, we would convert to feet per second. The answers we got indicate the asker was not exactly precise with the notation.

OpenStudy (anonymous):

we are doing conversions and my written problem was written like this...........30 m/h = _____________ft/s. I thought it was meters too.

OpenStudy (anonymous):

I got 2 answers with 44 ft/s......so I believe that answer is right.

OpenStudy (anonymous):

mph !=m/h

OpenStudy (anonymous):

I have a hard time setting up the conversions.....I get confused.....thank you though.

OpenStudy (anonymous):

In a practical sense, it doesn't seem that if something is moving really slow (30 meters per hour), you would want to know how fast that is in feet per second.

OpenStudy (anonymous):

A piece of metal is 3 3/4 ft x 4 2/3 ft. What is the area in square inches?

OpenStudy (turingtest):

@etina02 please post each question separately thanks